The Jackson-Timberlake Wardrobe Co. just paid a dividend of $1.56 per share on its stock. The dividends are expected to grow at a constant rate of 8 percent per year indefinitely.

  

a. If investors require a return of 13 percent on the company's stock, what is the current price?

  

b. What will the price be in 11 years?
Answer #1

a.Current price=D1/(Required return-Growth rate)

=(1.56*1.08)/(0.13-0.08)

=$33.696

b.P11=Current price*(1+Growth rate)^11

=33.696*(1.08)^11

=$78.57(Approx).